# Game Theory and Multi-Agent Reinforcement Learning for Zonal Ancillary Markets

### v3

@@ Abstract
We characterize zonal ancillary market coupling relying on noncooperative game theory. To that purpose, we formulate the ancillary market as a multi-leader single follower bilevel problem, that we subsequently cast as a generalized Nash game with side constraints and nonconvex feasibility sets. We determine conditions for equilibrium existence and show that the game has a generalized potential game structure. To compute market equilibrium, we rely on two exact approaches: an integrated optimization approach and Gauss-Seidel best-response, that we compare against multi-agent deep reinforcement learning. On real data from Germany and Austria, simulations indicate that multi-agent deep reinforcement learning achieves the smallest convergence rate but requires pretraining, while best-response is the slowest. On the economics side, multi-agent deep reinforcement learning results in smaller market costs compared to the exact methods, but at the cost of higher variability in the profit allocation among stakeholders. Further, stronger coupling between zones tends to reduce costs for larger zones.
Index Terms:
Ancillary Markets, Game Theory, MARL.
@@ I Introduction
A characteristic feature of electricity is the impossibility to store it in large proportions. Hence, it must be produced exactly when it is consumed, which raises a need for a real-time monitoring of the balance between production and consumption to guarantee the stability of the power grid. When imbalances occur, mechanisms are automatically triggered to restore the nominal frequency of the electric system: a change in power production is made by some production units, which are then said to provide ancillary services, in exchange of a monetary contribution. In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand. From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires some information about the market participants. This knowledge, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect. Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries. Market participants might be willing to investigate the impact of this shared market structure and to see if it stabilizes their revenues over time. The main challenges to address are the characterization of the market equilibria and the understanding of their evolution when market zones get interconnected: from the perspective of the market operator, it is important to gain insight into the behaviors of the stakeholders and to gauge the opportunity to enforce market integration over multiple countries if this can, e.g., mitigate price peak events in one country thanks to interconnection. In this article, we use noncooperative game theory to model ancillary services markets. Market equilibria are computed relying on reformulations of the game and multi-agent deep reinforcement learning (MARL). Our flexible framework allows us to assess the consequences of the interconnection of market zones by dynamically changing the coupling factors. The approach can be generalized to any number of zones.
@@
  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods, to compute an equilibrium ; a multi-agent deep reinforcement learning (MARL) approach, involving strategic learning and non-learning agents, that can be implemented in a decentralized fashion, reaching a stationary solution;
@@
  • we compare the three algorithm performance on real data from the Germany-Austria market, computing social cost and fairness indicators. Finally, we test multiple coupling settings for the market, observing a decrease in the market operators’ costs with higher couplings.
@@
This paper is organized as follows: in Section I-A, we introduce the relevant literature; in Section II, we formulate the ancillary service market as a noncooperative game. In Section III, we derive conditions for existence of market equilibria, that can be subsequently computed relying on the algorithms introduced in Section IV. Finally, in Section V, we implement the algorithms on real data from Austria and Germany, that we compare numerically using performance measures and economic metrics, concluding in Section VI.
I-A Related Work
There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers.
In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products. Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory. In [4] a two-stage electricity market design is considered, where the first stage is the day-ahead market and the second is the real time one. The authors obtain proof of equilibrium existence for the model and suggest a market mechanism to reach it. On ancillary markets, the two parts work of [5, 6], first discusses the design of an ancillary service focused on primary frequency response, and then simulates this design on different test cases. Another example exploring the impact of renewable based generators is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems. We conclude this part on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters.
Regarding studies of strategic bidding in electricity market, we start by mentioning [9], where the authors focus on the spot market, obtaining a Bayesian Nash equilibrium for the problem, computed in a centralized way and also considering the effect of wind based generation. In [4] a two-stage electricity market Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [10], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria. In [11] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources. The authors in [12] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program in order to solve it. This solution is computed from an external point of view, considering all producers at once. In [13], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints.
The use of reinforcement learning (RL) as a technique to solve complex problems related to power systems and electricity markets has become more common in the last years, even though applying to real world scenarios remains complicated due to the few theoretical results. One interesting work in this direction is [14], where the authors consider constraints to obtain guarantees for the application of an RL algorithm to voltage control. Regarding RL for strategic bidding in the electricity markets we mention [15], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers. Focusing specifically ancillary markets, the application of learning based algorithm is not yet widespread, nonetheless, there are some examples. In [16], the authors focus on ancillary markets with hydro-based producers. The simulation they propose uses discrete inputs for the algorithm and discretized bids for the learning producers, having market power. The results are not definitive, but the learning agents show sign of strategic behavior. A more advanced approach can be found in [17], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings. The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
@@ I-A Related Work
We denote $\mathbb{I}$ the indicator function. Given $\mathcal{A}\subset\mathbb{R}^{n}$, $|\mathcal{A}|$ refers to the cardinality of the set. Let $x_{n}\in\mathbb{R}^{m_{n}}$, with $m_{n}\in\mathbb{N}^{\star}$, we define $\text{col}((x_{n})_{n})$ as the stack of the $(x_{n})_{n}$ vectors. Given a vector of decision variables $x:=(x)_{n\in\mathcal{N}}$, we define $x_{-n}:=(x_{m})_{m\in\mathcal{N}\setminus\{n\}}$. When we write $\nabla^{2}f(x^{\star})$, with $x^{\star}\in\mathbb{R}^{d}$, $d\in\mathbb{R}^{\star}$, we consider the Hessian matrix of the function $f$ evaluated in $x^{\star}$, where the element $(i,j)$ refers to $\partial_{x_{i},x_{j}}f(x^{\star}),1\leq i,j\leq d$.
@@
$\mathcal{N}(z)$ Subset of producers in zone $z$
@@ II An Adversarial Game Formulation
The ancillary services market can be formulated as a multi-leader single follower bilevel problem. At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing. At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints. Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero sum game.
We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones. Each producer $n\in\mathcal{N}$ can bid up to $K_{n}\in\mathbb{N}^{\star}$ bids. Thus, the maximum number of bids is $B:=\sum_{n\in\mathcal{N}}K_{n}$. Further, each zone $z\in\mathcal{Z}$ has three input parameters: the total requested demand $D_{z}\geq 0$, the export limit $E_{z}\geq 0$ and the core portion $C_{z}\geq 0$. For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$. Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively. The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone. A summary of the definitions is in Tab. I.
@@ II-A Producers
$J_{n}(y_{n},x_{n}):=\sum_{k=1}^{K_{n}}\pi_{n,k}\Delta_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right).$ (2)
@@ II-B Market Operator
The market operator deals with four sets of constraints that depend on the capacities $\Delta_{n,k},\forall k\in[1,K_{n}],n\in\mathcal{N}$ submitted by the producers:
@@
Eq. (3) ensures that the total demand in each zone is met, Eq. (6) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq. (5) gives a cap to the amount of energy that can be exported. Finally, Eq. (4) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized. Constraints (3)-(6) can be compactly written under the form
@@
and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$. The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form:
@@
The market operator determines the fractions of bids to activate in each zone $(x_{n,k}^{z})_{n,k,z}$, under constraints (3)-(6):
@@
We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different zonal constraints.
@@ II-C Market Clearing and KKT Reformulation
We formulate the ancillary market clearing problem as a pessimistic multi-leader single follower bilevel optimization problem:
$\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(y_{n},x_{n}),$ (9a)
@@
The Lagrangian function associated with the market clearing defined in problem (8) takes the form:
@@
where $\Lambda\in\mathbb{R}_{+}^{3Z+B}$ refers to the dual variables associated with constraints (3)-(6). The analytical expression of the Lagrangian function in (10) can be found in Appendix A.
@@
    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
@@
Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution. Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0$. In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$. Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$. ∎
When the lower-level problem (8) satisfies Lemma 1, we can replace it by its KKTs in (9). Hence, each producer has to anticipate the decision $x$ of the market operator, meaning that in the single level the decision variables for producer $n$ will be $y_{n}\in\mathcal{Y}_{n}$ and $x_{n}\in\mathcal{X}_{n}(y)$. Before writing the full reformulation, we introduce tracking variables for every $z\in\mathcal{Z}$:
@@
The single-level problem for agent $n$ is
@@
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\tilde{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $3K_{n}+Z(K_{n}+1)+2$. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
Using the previously introduced compact notation with the explicit objective function, we get $J_{n}(w_{n},w_{-n})$. Thus, the $N$ optimization problems in Eqs. (12) give rise to a generalized Nash game $\Gamma$ that we denote in strategic form $\Gamma:=(\mathcal{N},\prod_{n}\mathcal{W}_{n},(J_{n})_{n})$:

### v4

@@ Abstract
We characterize zonal ancillary market coupling relying on noncooperative game theory. We formulate the ancillary market as a multi-leader single follower Stackelberg game, that we subsequently cast as a generalized Nash game with side constraints and nonconvex feasible sets. We determine conditions for the existence of an equilibrium and show that the game has a generalized potential structure. To compute market equilibrium, we rely on two exact approaches: an integrated optimization reformulation involving complementarity constraints, that can be solved using a nonlinear optimization solver, and Gauss-Seidel best-response. We compare these exact methods against multi-agent deep reinforcement learning (MARL). Simulations on real data from Germany and Austria ancillary markets show that MARL better capture intricate interactions between market agents under varying information-sharing scenarios, achieving faster learning rates to reach a stationary solution, than both exact approaches. Further, MARL results in smaller market costs compared to the exact methods, at the cost of higher variability in the profit allocation among stakeholders. Finally, on the policy side, we show numerically that tighter coupling between adjacent supply zones reduces the average market cost of the largest zone.
@@ I Introduction
A characteristic feature of electricity is the impossibility to store it in large proportions. Hence, it must be produced exactly when it is consumed, which raises a need for a real-time monitoring of the balance between production and consumption to guarantee the stability of the power grid. When imbalances occur, mechanisms are automatically triggered to restore the nominal frequency of the electric system: a change in power production is made by some production units, which are then said to provide ancillary services, in exchange of a monetary contribution. In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand. From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires information about the other market participants. This information, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect. Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries. Market participants might be willing to investigate the impact of this shared market structure and to see if it stabilizes their revenues over time. The main challenges to address are the characterization of the market equilibria and the understanding of their evolution when market zones get interconnected: from the perspective of the market operator, it is important to gain insight into the behaviors of the stakeholders and to gauge the opportunity to enforce market integration over multiple countries if this can, e.g., mitigate price peak events in one country thanks to interconnection. In this article, we use noncooperative game theory to model ancillary services markets. Market equilibria are computed relying on reformulations of the game and multi-agent deep reinforcement learning (MARL). Our flexible framework allows us to assess the consequences of the interconnection of market zones by dynamically changing the coupling factors. The approach can be generalized to any number of zones.
I-A Related Work
There are two main directions for research on ancillary (and zonal) market design: a) working on the design of the market itself, studying different formulations and their features [lecadre2019]; b) working on the strategic bidding aspect [morri2024], focusing more on the point of view of the energy producers. In [oren_design_2001], Oren discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products. Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory. A new market design is proposed by Ela et al. in [ela_effective_2012], taking into account the uncertainty resulting from the massive penetration of wind power in the grid. In the same vein, [ela_market_2014, ela_market_2014_2] discuss the design of an ancillary services market focused on primary frequency response, and then simulate this design under different scenarios. Another example exploring the impact of renewable based generators is [merten2020bidding], which tackles the optimization of bidding strategies for ancillary markets involving battery energy storage systems. We conclude the discussion on the formulation and features of ancillary services markets with [oureilidis_ancillary_2020], which explores the pros. and cons. of ancillary services market implementation at the distribution grid level, considering frequency response, voltage power regulation and other grid related parameters. Regarding the computation of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [morinec2008optimal], where the authors model the market as a two-player noncooperative game, analyzing the strategies of the two producers, and solving the game to obtain different market equilibria. In [meyn_ancillary_2014], a control based decentralized solution is proposed to create ancillary services helping with power imbalances caused by renewable generation sources. Sarfati and Holmberg in [sarfati_simulation_2020] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program that they subsequently solve relying on a nonlinear optimization solver. This solution is computed from an external point of view, considering all producers at once. In [jay2020game], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints. Finally, [deng_distributed_2024] focuses on a game-theoretic model with nonlinear convex feasible sets, which in turn is solved by simulation relying on a linear dynamic model. The authors then analyze theoretically the existence of a variational equilibrium and propose a distributed algorithm to compute it. The use of learning algorithm for strategic bidding has become more common in the last years. In [ye_deep_2020], a day-ahead electricity market bidding problem with multiple strategic producers is formulated as a Markov game with unobservable information, which is solved relying on a model-free and data-driven approach. Regarding ancillary markets, the application of learning based algorithm is not particularly widespread, nonetheless, there are some examples. In [abgottspon_strategic_2013], Abgottspon and Andersson focus on ancillary markets with hydro-based producers. Their simulation approach uses discrete inputs and discretized bids for the learning producers. The results are not definitive, but the learning agents show sign of strategic behavior. A more advanced approach can be found in [du_approximating_2021], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings. The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
I-B Contributions and Organization
@@
  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods; a multi-agent deep reinforcement learning (MARL) approach, that can be implemented in a decentralized fashion, reaching a stationary solution;
@@
  • we compare the three algorithm performance on real data from the Germany-Austria market, comparing social cost, producers’ profit variability and computation times. Finally, comparing multiple coupling scenarios, we observe a decrease in the largest zone market cost with tighter couplings.
@@
This paper is organized as follows: in Section II, we formulate the ancillary services market as a noncooperative game, in Section III, we derive conditions for existence of market equilibria, that can be subsequently computed relying on the algorithms introduced in Section IV. Finally, in Section V, we implement the algorithms on real data from Austria and Germany, that we compare numerically using key performance metrics. We conclude in Section VI.
@@
Given $\mathcal{A}\subset\mathbb{R}^{n}$, $|\mathcal{A}|$ refers to the cardinality of the set. Let $x_{n}\in\mathbb{R}^{m_{n}}$, with $m_{n}\in\mathbb{N}^{\star}$, we define $\text{col}((x_{n})_{n})$ as the stack of the $(x_{n})_{n}$ vectors. Given a vector of decision variables $x:=(x_{n})_{n\in\mathcal{N}}$, we define $x_{-n}:=(x_{m})_{m\in\mathcal{N}\setminus\{n\}}$ as the concatenation of all the players’ decision variables except $n$. When we write $\nabla^{2}f(x^{\star})$, with $x^{\star}\in\mathbb{R}^{d}$, $d\in\mathbb{R}^{\star}$, we consider the Hessian matrix of the function $f$ evaluated in $x^{\star}$, where the element $(i,j)$ refers to $\partial_{x_{i},x_{j}}f(x^{\star}),\,\forall i,j\in\{1,\ldots,d\}$.
@@
$\mathcal{N}(z)$ Subset of producers in $z$
@@ II An Adversarial Game Formulation
The ancillary services market can be formulated as a multi-leader single follower Stackelberg game. At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing. At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints. Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero-sum game.
We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones. Each producer $n\in\mathcal{N}$ can bid up to $K_{n}\in\mathbb{N}^{\star}$ bids. Thus, the maximum number of bids is $B:=\sum_{n\in\mathcal{N}}K_{n}$. Further, each zone $z\in\mathcal{Z}$ has three input parameters: the total requested demand $D_{z}\geq 0$, the export limit $E_{z}\geq 0$ and the core portion $C_{z}\geq 0$. For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$. Each producer $n$’s decision variable is a finite sequence of bids
$y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big((\Delta_{n,l},\pi_{n,l})_{l=1,\ldots,K_{n}}\Big),$ (1)
where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively. We denote by $y:=(y_{n})_{n}$ the producers’ strategy profile.
The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone. A summary of the definitions is displayed in Tab. I.
@@ II-A Producers
$J_{n}(x_{n},y_{n}):=\sum_{k=1}^{K_{n}}\pi_{n,k}\Delta_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right).$ (3)
@@ II-B Market Operator
The market operator deals with four sets of constraints that depend on the capacities $\Delta_{n,k},\forall k\in[1,K_{n}],n\in\mathcal{N}$ submitted by the producers
@@
Eq. (4) ensures that the total demand in each zone is met, Eq. (7) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq. (6) gives a cap to the amount of energy that can be exported. Finally, Eq. (5) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized. To highlight the mathematical structure of the problem, constraints (4)-(7) can be compactly written under the form
@@
where $y:=(y_{n})_{n}$, $y_{n}$ is defined in Eq. (1), and $\Phi:[0,1]^{B\times Z}\times\prod_{n\in\mathcal{N}}[[\underline{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}]\to\mathbb{R}^{3Z+N}$. We denote
$\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}\,|\,\Phi(x,y)\leq 0\},$
the feasible set of the market operator which depends on the profile of strategies of all the producers $y:=(y_{n})_{n}$. The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form
@@
The market operator determines the fractions of bids to activate in each zone $(x_{n,k}^{z})_{n,k,z}$, under constraints (4)-(7)
@@
We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different constraints.
Zone AZone BZonal Demand AZonal Demand BExportCore PortionCore PortionP1P2P3P4P5
@@ II-C Market Clearing and KKT Reformulation
We formulate the ancillary services market clearing problem as a pessimistic multi-leader single follower Stackelberg game
$\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(x_{n},y_{n}),$ (10a)
@@
The Lagrangian function associated with the market clearing defined in problem (9) takes the form
@@
where $\Lambda\in\mathbb{R}_{+}^{3Z+B}$ refers to the dual variables associated with constraints (4)-(7).
@@
    and
    $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
@@
Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution. Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e. $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0.$
In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus
$\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}.$
Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}.$
When the lower-level problem (9) satisfies Lemma 1, we can replace it by its KKTs in (10). Hence, each producer has to anticipate the decision $x$ of the market operator, meaning that in the single level the decision variables for producer $n$ will be $y_{n}\in\mathcal{Y}_{n}$ and $x_{n}\in\mathcal{X}_{n}(y)$ with $y=(y_{n})_{n}$. Before writing the full reformulation, we introduce tracking variables for every zone $z\in\mathcal{Z}$
@@
The single-level problem for agent $n$ takes the form
@@
We now aim to highlight the structure of the problem. We first note that, given the dependence of the feasible set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, $x_{n}(y)$. This allows to write the producer’s objective function as an explicit function of all producers’ decision variables, i.e., $J_{n}(y_{n},y_{-n})$.
However, in this reformulation, the dual variables are decision variables of the producers. To simplify, we need an additional change of variable. We set
$w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n},$
where $\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\underline{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$ is a polyhedron of dimension $3K_{n}+Z(K_{n}+1)+2$.
The set of the $l_{n}$ local constraints are represented by the function $h_{n}:\mathcal{W}_{n}\rightarrow\mathbb{R}^{l_{n}}$. Further, we introduce the local feasible set for each player $n$
$\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}.$
Similarly, the $r$ shared constraints are formulated in a compact form as $G:\Omega(w)\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}(w_{-n}):=\{w_{n}\in\widetilde{\mathcal{W}}_{n}\,|\,G(w_{n},w_{-n})\leq 0\}$ and $\Omega(w):=\prod_{n=1}^{N}\Omega_{n}(w_{-n})$. We check that $h_{n}$ and $G$ are continuously differentiable mappings.
Relying on these notations, the $N$-producer coupled optimization problems, derived from Eqs. (13), give rise to a nonconvex game with side constraints [pang_nonconvex_2011]
$\Gamma:=\Big(\mathcal{N},\Omega(\cdot),\left(J_{n}(\cdot)\right)_{n}\Big),$
where producers solve the parametrized optimization problems
@@
Pang and Scutari in [pang_nonconvex_2011] study the properties of problem (14), providing different sets of assumptions to be used in order to obtain existence guarantees for specific equilibria. In the next section, we rely on a result guaranteeing the existence of a Nash equilibrium for a relaxed version of the problem. Further, under mild assumptions on the model parameters, we show the equivalence between the set of solutions of the relaxed version and the generalized Nash game $\Gamma$. Refinements of generalized Nash equilibria (GNEs) solutions of $\Gamma$ requiring the uniform evaluation of the Lagrange multipliers associated with the coupling constraints across the producers, are called variational equilibria (v-GNEs) [harker1991]. Interestingly, the generalized potential structure of $\Gamma$, guarantees the economic efficiency of v-GNEs. However, reaching v-GNEs requires some implicit coordination between the market participants which may be difficult to enforce. This will be discussed further in Section V.